WHAT IS THE THE ANSWER 5b to the power 3. B=2
Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

5b^3
b=2
5*2^3
2^3=8
5*8=40

Answer by checkley77(12844)

5b to the power 3. B=2
5*2^3
5*8=40 ans.
or:
(5*2)^3
10^3=1,000 ans.